# Evaluation Guidelines for Empirical Studies in Software Engineering involving LLMs

### v1

@@ 4.7.2. Example(s)
According to Hou et al., main problems types for LLMs are classification, recommendation and generation problems. Each of these problem types requires a different set of metrics. Hu et al. conducted a comprehensive structured literature review on LLM benchmarks related to software engineering tasks. They analyzed 191 benchmarks and categorized them according to the specific task they address, making the paper a valuable resource for identifying existing metrics. Metrics used to assess generation tasks include BLEU, pass@k, Accuracy, Accuracy@k, and Exact Match. The most common metric for recommendation tasks is Mean Reciprocal Rank. For classification tasks, classical machine learning metrics such as Precision, Recall, F1-score, and Accuracy are often reported.
@@ 4.7.4. Challenges
Many closed-source models, such as those released by OpenAI, achieve exceptional performance on certain tasks but lack transparency and reproducibility (Li et al., 2024b; Du et al., 2023; Zhuo et al., 2024). Benchmark leaderboards, particularly for code generation, are led by close-sourced models (Du et al., 2023; Zhuo et al., 2024). While researchers should compare performance against these models, they must consider that providers might discontinue them or apply undisclosed pre- or post-processing beyond the researchers’ control (see Use an Open LLM as a Baseline).
@@ 4.8.1. Recommendations
A cornerstone of open science is the ability to reproduce research results. Although the inherent non-deterministic nature of LLM is a strength in many use cases, its impact on reproducibility is a challenge. To enable reproducibility, researchers should disclose a replication package for their study. They should perform multiple repetitions of their experiments (see Report Suitable Baselines, Benchmarks, and Metrics) to account for non-deterministic outputs. In some cases, researchers may reduce the output variability by setting the temperature to a value close to 0 and setting a fixed seed value. However, configuring a lower temperature can negatively impact task performance, and not all models allow the configuration of seed values. Besides the inherent non-determinism, the behavior of an LLM depends on many external factors such as prompt variations and model evolution. To ensure reproducibility, researchers should follow all previous guidelines and further discuss the generalization challenges outlined below.
Although the topic of generalizability is not new, it has gained new relevance with increasing interest in LLMs. In LLM-based studies, generalizability boils down to two main concerns: (1) First, are the results specific to an LLM, or can they be achieved with other LLMs as well? (2) Second, will these results still be valid in the future? If generalizability to other LLMs is not in the scope of the research, this must be clearly explained in the paper or, if generalizability is in scope, researchers must compare their results or subsets of the results (e.g., due to computational cost) with other LLMs that are, e.g., similar in size, to assess the generalizability of their findings (see also Section Use an Open LLM as a Baseline). Multiple studies (e.g., (Chen et al., 2023; Li et al., 2024a)) found that the performance of proprietary LLMs (e.g., GPT) decreased over time for certain tasks using the same model version (e.g. GPT-4o). Reporting the model version and configuration is not sufficient in such cases. To date, the only way to mitigate this limitation is the usage of an open LLM having a versioning schema and being archiving (see Use an Open LLM as a Baseline). Hence, researchers should employ open LLMs to establish a reproducible baseline and, if the use of an open LLM is not possible, researchers should test and report their results over an extended period of time as a proxy of the stability of the results over time.
@@
Conducting studies with LLMs is a resource-intensive process. For self-hosted LLMs, the respective hardware must be provided, and for managed LLMs, the service costs must be considered. The challenge becomes more pronounced as LLMs grow larger, research architectures become more complex, and experiments become more computationally expensive. For example, multiple repetitions to assess model or tool performance (see Report Suitable Baselines, Benchmarks, and Metrics) multiply the cost and impact scalability. Consequently, resource-intensive research remains predominantly the domain of private companies or well-funded research institutions, hindering researchers with limited resources in reproducing, replicating, or extending study results. Hence, for transparency reasons, researchers should report the cost associated with executing an LLM-based study. If the study used self-hosted LLMs, researchers should report the specific hardware used. If the study used managed LLMs, the service cost should be reported. To ensure the validity and reproducibility of the research results, researchers must provide the LLM outputs as evidence (see Section Report Prompts, their Development, and Interaction Logs). Depending on the architecture (see Section Report Tool Architecture beyond Models), these outputs should be reported on different architectural levels (e.g., outputs of individual LLMs in multi-agent systems). Additionally, researchers should include a subset of the employed validation dataset, selected using an accepted sampling strategy, to allow partial replication of the results.
@@
Although metrics such as BLEU or ROUGE are commonly used to evaluate the performance of LLMs (see Section Report Suitable Baselines, Benchmarks, and Metrics), they may not capture other relevant SE-specific aspects such as functional correctness or runtime performance of automatically generated source code (Liu et al., 2023). Given the high resource demand of LLMs, in addition to traditional metrics such as accuracy, precision, and recall or more contemporary metrics such as pass@k, resource consumption has to become a key indicator of performance. While research has predominantly focused on energy consumption during the early phases of building LLMs (e.g., data center manufacturing, data acquisition, training), inference, that is LLM usage, often becomes similarly or even more resource-intensive (de Vries, 2023; Wu et al., 2022; Fu et al., 2024; Jiang et al., 2024; Mitu and Mitu, 2024). Hence, researchers should aim for lower resource consumption on the model side. This can be achieved by selecting smaller (e.g., GPT-4o-mini instead of GPT-4o) or newer models or by employing techniques such as model pruning, quantization, or knowledge distillation (Mitu and Mitu, 2024). Researchers may further reduce resource consumption by restricting the number of queries, input tokens, or output tokens (Mitu and Mitu, 2024), by using different prompt engineering techniques (e.g., zero-shot prompts seem to emit less CO2 than chain-of-thought prompts), or by carefully sampling smaller datasets for fine-tuning and evaluation instead of using large datasets. Reducing resource consumption involves a trade-off with our recommendation to perform multiple experimental runs to account for non-determinism, as suggested in Report Suitable Baselines, Benchmarks, and Metrics. To report the environmental impact of a study, researchers should use software such as CodeCarbon or Experiment Impact Tracker to track and quantify the carbon footprint of the study or report an estimate of the carbon footprint through tools such as MLCO2 Impact. They should detail the LLM version and configuration as described in Report Model Version, Configuration, and Customizations, state the hardware or hosting provider of the model as described in Report Tool Architecture beyond Models and report the total number of requests, accumulated input and output tokens. Researchers must justify why LLMs were chosen over existing approaches and discuss the achieved performance in relation to their potentially higher resource consumption.
@@ 4.8.2. Example(s)
Bias can occur in LLM training datasets, resulting in various types of discrimination. Gallegos et al. propose metrics to quantify biases in various tasks (e.g., text generation, classification, question answering) (Gallegos et al., 2023).
@@ 4.8.4. Challenges
With commercial LLMs evolving over time, the generalizability of results to future model versions is uncertain. Employing open LLMs as a baseline can mitigate this limitation, but may not always be feasible due to computational cost. Most LLM providers do not publicly offer information about their training data, impeding the assessment of inter-dataset duplication effects. Consistently keeping track of and reporting the costs involved in a research endeavor is challenging. Building a coherent replication package that includes LLM outputs and samples for partial replicability requires additional effort and resources. Finding the right metrics to evaluate the performance of LLMs in SE for specific tasks is challenging. Defining all requirements beforehand to ensure the use of suitable metrics can be challenging, especially in exploratory research. Our Section Report Suitable Baselines, Benchmarks, and Metrics and its references can serve as a starting point. Ensuring compliance across jurisdictions is difficult with different regions having different regulations and requirements (e.g., GDPR and the AI Act in the EU, CCPA in California). Selecting datasets and models with fewer biases is challenging, as bias are often unknown. Measuring or estimating the environmental impact of a study is challenging and might not always be feasible. Especially in exploratory research, the impact is hard to estimate in advance, making it difficult to justify the usage of LLMs over other approaches.
@@ 5. Conclusion
This paper organizes the landscape of empirical studies involving LLMs in software engineering along seven study types and distills eight guidelines aimed at improving their validity, transparency, and reproducibility. Our guidelines recommend researchers to (1) explicitly declare when and how LLMs are used; (2) report model versions, configuration, and fine-tuning details; (3) describe the complete tool architecture beyond the model; (4) release prompts, their development, and, where possible, interaction logs; (5) validate LLM outputs with humans; (6) include an open LLM as a baseline; (7) select appropriate baselines, benchmarks, and metrics; and (8) report study limitations and mitigations, including costs, potential biases, and environmental impact. Adopting these practices will not eliminate LLMs’ non-determinism, but it will make claims auditable and results easier to replicate and compare across studies. By consistently applying our recommendations, the community can produce more reliable evidence on what LLM can and cannot deliver for software engineering research and practice. Our guidelines complement existing empirical software engineering guidance while addressing LLM-specific challenges. Because models and tools evolve, we maintain the guidelines as a living resource and invite the community to refine and extend them: llm-guidelines.org.

### v2

@@ 5.7.2. Example(s)
According to Hou et al., main problem types for LLMs are classification, recommendation and generation problems. Each of these problem types requires a different set of metrics. Hu et al. conducted a comprehensive structured literature review on LLM benchmarks related to software engineering tasks. They analyzed 191 benchmarks and categorized them according to the specific task they address, making the paper a valuable resource for identifying existing metrics. Metrics used to assess generation tasks include BLEU, pass@k, Accuracy, Accuracy@k, and Exact Match. The most common metric for recommendation tasks is Mean Reciprocal Rank. For classification tasks, classical machine learning metrics such as Precision, Recall, F1-score, and Accuracy are often reported.
@@ 5.7.4. Challenges
Many closed-source models, such as those released by OpenAI, achieve exceptional performance on certain tasks but lack transparency and reproducibility (Li et al., 2024b; Du et al., 2023; Zhuo et al., 2024). Benchmark leaderboards, particularly for code generation, are led by closed-source models (Du et al., 2023; Zhuo et al., 2024). While researchers should compare performance against these models, they must consider that providers might discontinue them or apply undisclosed pre- or post-processing beyond the researchers’ control (see Use an Open LLM as a Baseline).
@@ 5.8.1. Recommendations
A cornerstone of open science is the ability to reproduce research results. Although the inherent non-deterministic nature of LLM is a strength in many use cases, its impact on reproducibility is problematic. To enable reproducibility, researchers should disclose a replication package for their study. They should perform multiple repetitions of their experiments (see Use Suitable Baselines, Benchmarks, and Metrics) to account for non-deterministic outputs. In some cases, researchers may reduce the output variability by setting the temperature to a value close to 0 and setting a fixed seed value. However, configuring a lower temperature can negatively impact task performance, and not all models allow the configuration of seed values. Besides the inherent non-determinism, the behavior of an LLM depends on many external factors such as prompt variations and model evolution. To ensure reproducibility, researchers should follow all previous guidelines and further discuss the generalization challenges outlined below.
Although the topic of generalizability is not new, it has gained new relevance with increasing interest in LLMs. In LLM-based studies, generalizability boils down to two main concerns: (1) First, are the results specific to an LLM, or can they be achieved with other LLMs as well? (2) Second, will these results still be valid in the future? If generalizability to other LLMs is not in the scope of the research, this must be clearly explained in the paper or, if generalizability is in scope, researchers must compare their results or subsets of the results (e.g., due to computational cost) with other LLMs that are, e.g., similar in size, to assess the generalizability of their findings (see also Section Use an Open LLM as a Baseline). Multiple studies (e.g., (Chen et al., 2023; Li et al., 2024a)) found that the performance of proprietary LLMs (e.g., GPT) decreased over time for certain tasks using the same model version (e.g. GPT-4o). Reporting the model version and configuration is not sufficient in such cases. To date, the only way to mitigate this limitation is the usage of an open LLM having a versioning schema and archiving (see Use an Open LLM as a Baseline). Hence, researchers should employ open LLMs to establish a reproducible baseline and, if the use of an open LLM is not possible, researchers should test and report their results over an extended period of time as a proxy of the stability of the results over time.
@@
Conducting studies with LLMs is a resource-intensive process. For self-hosted LLMs, the respective hardware must be provided, and for managed LLMs, the service costs must be considered. The challenge becomes more pronounced as LLMs grow larger, research architectures become more complex, and experiments become more computationally expensive. For example, multiple repetitions to assess model or tool performance (see Use Suitable Baselines, Benchmarks, and Metrics) multiply the cost and impact scalability. Consequently, resource-intensive research remains predominantly the domain of private companies or well-funded research institutions, hindering researchers with limited resources in reproducing, replicating, or extending study results. Hence, for transparency reasons, researchers should report the cost associated with executing an LLM-based study. If the study used self-hosted LLMs, researchers should report the specific hardware used. If the study used managed LLMs, the service cost should be reported. To ensure the validity and reproducibility of the research results, researchers must provide the LLM outputs as evidence (see Section Report Prompts, their Development, and Interaction Logs). Depending on the architecture (see Section Report Tool Architecture beyond Models), these outputs should be reported on different architectural levels (e.g., outputs of individual LLMs in multi-agent systems). Additionally, researchers should include a subset of the employed validation dataset, selected using an accepted sampling strategy, to allow partial replication of the results.
@@
Although metrics such as BLEU or ROUGE are commonly used to evaluate the performance of LLMs (see Section Use Suitable Baselines, Benchmarks, and Metrics), they may not capture other relevant SE-specific aspects such as functional correctness or runtime performance of automatically generated source code (Liu et al., 2023). Given the high resource demand of LLMs, in addition to traditional metrics such as accuracy, precision, and recall or more contemporary metrics such as pass@k, resource consumption has to become a key indicator of performance. While research has predominantly focused on energy consumption during the early phases of building LLMs (e.g., data center manufacturing, data acquisition, training), inference, that is LLM usage, often becomes similarly or even more resource-intensive (de Vries, 2023; Wu et al., 2022; Fu et al., 2024; Jiang et al., 2024; Mitu and Mitu, 2024). Hence, researchers should aim for lower resource consumption on the model side. This can be achieved by selecting smaller (e.g., GPT-4o-mini instead of GPT-4o) or newer models or by employing techniques such as model pruning, quantization, or knowledge distillation (Mitu and Mitu, 2024). Researchers may further reduce resource consumption by restricting the number of queries, input tokens, or output tokens (Mitu and Mitu, 2024), by using different prompt engineering techniques (e.g., zero-shot prompts seem to emit less CO2 than chain-of-thought prompts), or by carefully sampling smaller datasets for fine-tuning and evaluation instead of using large datasets. Reducing resource consumption involves a trade-off with our recommendation to perform multiple experimental runs to account for non-determinism, as suggested in Use Suitable Baselines, Benchmarks, and Metrics. To report the environmental impact of a study, researchers should use software such as CodeCarbon or Experiment Impact Tracker to track and quantify the carbon footprint of the study or report an estimate of the carbon footprint through tools such as MLCO2 Impact. They should detail the LLM version and configuration as described in Report Model Version, Configuration, and Customizations, state the hardware or hosting provider of the model as described in Report Tool Architecture beyond Models and report the total number of requests, accumulated input and output tokens. Researchers must justify why LLMs were chosen over existing approaches and discuss the achieved performance in relation to their potentially higher resource consumption.
@@ 5.8.2. Example(s)
Finally, bias can occur in LLM training datasets, resulting in various types of discrimination. Gallegos et al. propose metrics to quantify biases in various tasks (e.g., text generation, classification, question answering) (Gallegos et al., 2023).
@@ 5.8.4. Challenges
With commercial LLMs evolving over time, the generalizability of results to future model versions is uncertain. Employing open LLMs as a baseline can mitigate this limitation, but may not always be feasible due to computational cost. Most LLM providers do not publicly offer information about their training data, impeding the assessment of inter-dataset duplication effects. Consistently keeping track of and reporting the costs involved in a research endeavor is challenging. Building a coherent replication package that includes LLM outputs and samples for partial replicability requires additional effort and resources. Finding the right metrics to evaluate the performance of LLMs in SE for specific tasks is challenging. Defining all requirements beforehand to ensure the use of suitable metrics can be a challenge, especially in exploratory research. Our Section Use Suitable Baselines, Benchmarks, and Metrics and its references can serve as a starting point. Ensuring compliance across jurisdictions is difficult with different regions having different regulations and requirements (e.g., GDPR and the AI Act in the EU, CCPA in California). Selecting datasets and models with fewer biases is challenging, as biases are often unknown. Measuring or estimating the environmental impact of a study is challenging and might not always be feasible. Especially in exploratory research, the impact is hard to estimate in advance, making it difficult to justify the usage of LLMs over other approaches.
@@ 6. Conclusion
This paper organizes the landscape of empirical studies involving LLMs in software engineering along seven study types and distills eight guidelines aimed at improving their validity, transparency, and reproducibility. Our guidelines recommend that researchers (1) explicitly declare when and how LLMs are used; (2) report model versions, configuration, and fine-tuning details; (3) describe the complete tool architecture beyond the model; (4) release prompts, their development, and, where possible, interaction logs; (5) validate LLM outputs with humans; (6) include an open LLM as a baseline; (7) select appropriate baselines, benchmarks, and metrics; and (8) report study limitations and mitigations, including costs, potential biases, and environmental impact. Adopting these practices will not eliminate LLM non-determinism, but it will make claims auditable, and results easier to replicate and compare across studies. By consistently applying our recommendations, the community can produce more reliable evidence on what LLM can and cannot deliver for software engineering research and practice. Our guidelines complement existing empirical software engineering guidance while addressing LLM-specific challenges. Because models and tools evolve, we maintain the guidelines as a living resource and invite the community to refine and extend them: llm-guidelines.org.